State amplitude and period for each function y=3cos4 theta and y= -2 sin (3pie over 2) theta
Answer by lwsshak3(11628)

State amplitude and period for each function y=3cos4 theta and y= -2 sin (3pie over 2) theta
**
Equation for cos function:
y=Acos(Bx-C),A=amplitude, Period=2π/B, C/B=Phase shift
For given equation y=3cos4x:
Amplitude=3
B=4
Period=2π/B
=2π/4
x=π/2
phase-shift: none
..
Equation for sin function:
y=Asin(Bx-C),A=amplitude, Period=2π/B, C/B=Phase shift
For given equation y=-2sin(3πx/2)
Amplitude=2
B=3π/2
Period=2π/B
=2π/(3π/2)
x=4/3
phase-shift: none